OpenStudy (anonymous):

Tom can mow a lawn in 4 hours. It takes Perry 5 hours to mow the same lawn. How long would it take both of them, working together, to mow the same lawn?

jimthompson5910 (jim_thompson5910):

1/4 + 1/5 = 1/t 5/20 + 4/20 = 1/t 9/20 = 1/t 9t = 20*1 9t = 20 t = 20/9 So it will take them 20/9 hours, or 2 and 2/9 hours, to get the job done if they do it together.

OpenStudy (anonymous):

thank you SO much!!!

jimthompson5910 (jim_thompson5910):

np